# DWTNeRF: Boosting Few-shot Neural Radiance Fields via Discrete Wavelet Transform

### v2

@@ 13 Model methods
In previous sections, we established the advantages of DWTNeRF’s DW loss over two prevalent prior-based losses, while validating the design choices. In this section, we compared DWTNeRF’s cross-branch interactions against model methods as introduced by mi-MLP [46]: “Residual Connections” and “Element-wise Cross-branch Interactions”. The architectures of each method are provided in Figure 6. In general, model methods aim to boost few-shot NeRF by tweaking Vanilla MLP’s architecture.
@@
Network Structures Following CombiNeRF [2] and mi-MLP [46], DWTNeRF is modelled using separate MLP branches for densities and colors. The density branch has 2 MLP layers, and its inputs are encoded using hash encoding [22] with 16 resolution levels. The color branch has 3 MLP layers, and its inputs are encoded using 64 bases of the spherical harmonics [27].
@@
Residual Connections This provides a shorter path between the inputs and intermediate MLP layers, helping to reduce overfitting. For our implementation onto INGP, we introduced the residuals with the concatenation operator [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]:
@@
Element-wise Cross-branch Interactions This provides interactions between color and density branches, regularizing color predictions through position-aware relationships. We implemented this into INGP as follows:
@@
As presented in Table 4, all of the above methods degrade performance. Similar to some prior methods, model methods are not trivially transferable across Vanilla NeRF and INGP, and very possibly, 3DGS. This is due to INGP’s aggressive optimization dynamics, making it vulnerable to very slight architectural changes. Our method, based on multi-head attention and acting only on the inputs & outputs, bypasses this shortcoming while still capable of learning density-color interactions.

### v3

@@ 9 Model-based methods
In this section, we compared DWTNeRF’s cross-branch interactions against model-based methods as introduced by mi-MLP [46]: “Residual Connections” and “Element-wise Cross-branch Interactions”. The architectures of each method are provided in Figure 6. In general, model-based methods aim to boost few-shot NeRF by tweaking Vanilla MLP’s architecture.
@@
Network Structures. Following CombiNeRF [2] and mi-MLP [46], DWTNeRF is modelled using separate MLP branches for densities and colors. The density branch has 2 MLP layers, and its inputs are encoded using hash encoding [22] with 16 resolution levels. The color branch has 3 MLP layers, and its inputs are encoded using 64 bases of the spherical harmonics [27].
@@
Residual Connections. This provides a shorter path between the inputs and intermediate MLP layers, helping to reduce overfitting. For our implementation onto INGP, we introduced the residuals with the concatenation operator $\cdot$$\cdot$]:
@@
Element-wise Cross-branch Interactions. This provides interactions between color and density branches, regularizing color predictions through position-aware relationships. We implemented this into INGP as follows:
@@
As presented in Table 4, all of the above methods degrade performance. Similar to some prior-based methods, model-based methods are not trivially transferable across Vanilla NeRF and INGP, and very possibly, 3DGS. This is due to INGP’s aggressive optimization dynamics, making it vulnerable to very slight architectural changes. Our method, based on multi-head attention and acting only on the inputs & outputs, bypasses this shortcoming while still capable of learning density-color interactions.